### v2

# Youngjoon Lee1, Jinu Gong2, Joonhyuk Kang1 1School of Electrical Engineering, KAIST, South Korea

Abstract
Given sufficient data from multiple edge devices, federated learning (FL) enables training a shared model without transmitting private data to a central server. However, FL is generally vulnerable to Byzantine attacks from compromised edge devices, which can significantly degrade the model performance. In this paper, we propose a intuitive plugin that can be integrated into existing FL techniques to achieve Byzantine-Resilience. Key idea is to generate virtual data samples and evaluate model consistency scores across local updates to effectively filter out compromised edge devices. By utilizing this scoring mechanism before the aggregation phase, the proposed plugin enables existing FL techniques to become robust against Byzantine attacks while maintaining their original benefits. Numerical results on medical image classification task validate that plugging the proposed approach into representative FL algorithms, effectively achieves Byzantine resilience. Furthermore, the proposed plugin maintains the original convergence properties of the base FL algorithms when no Byzantine attacks are present.
Index Terms: e-health, federated learning, adversarial fault tolerance, virtual data, consistency scoring
I Introduction
Recent advances in deep learning have transformed healthcare by leveraging large-scale medical datasets [1, 2]. However, strict patient privacy regulations and HIPAA compliance requirements have created significant barriers to traditional centralized analysis of healthcare data [3]. Federated learning (FL) [4] has emerged as a promising framework for healthcare applications [5]. This approach enables collaborative training of medical AI models while preserving patient privacy through secure parameter sharing [6]. Clinical diagnosis and medical imaging analysis have shown significant improvements using this privacy-preserving paradigm [7]. These healthcare applications handle sensitive protected health information (PHI) that cannot be directly shared among medical institutions [8].
Figure 1: Comparison of conventional FL and plugin-enhanced FL under Byzantine attacks. In the conventional approach (top), malicious edge devices (devices 3 and 4) inject poisoned model updates that compromise the global model during aggregation. Our proposed plugin (bottom) effectively filters out these poisoned updates to maintain model integrity, leading to an optimal global model.
While FL offers privacy benefits, implementing it in real-world healthcare systems faces significant challenges [9]. Healthcare data collected from different medical institutions naturally exhibits data heterogeneity due to varying patient populations and clinical protocols [10, 11]. Moreover, security threats pose a critical concern in federated medical systems, as malicious participants can compromise the global model through Byzantine attacks with poisoned updates [12]. This vulnerability is particularly concerning in clinical applications where model failures could directly impact patient safety and care outcomes [13]. Therefore, developing robust defense mechanisms against adversarial attacks while maintaining model performance under heterogeneous conditions remains an important challenge.
Novel plugin-based architecture proposed as shown in Fig. 1, seamlessly integrates Byzantine resilience into existing FL techniques without compromising their original benefits. Our approach generates virtual data samples to evaluate model behavior patterns. Using these patterns, we compute consistency scores across local updates to identify and filter out malicious edge devices. Our modular approach maintains the core functionality of existing FL techniques. This intuitive plugin can be readily applied into various FL frameworks without requiring complete system redesign. Extensive experiments on medical imaging dataset show that our plugin improves robustness while maintaining high performance across FL algorithms. Main contributions of this paper are as follows:
  • We propose a novel plugin to embed Byzantine resilience into FL techniques without altering their core algorithms.
  • We propose a virtual data-driven scoring method to detect and filter compromised edge devices.
  • We show our plugin’s compatibility and validate its effectiveness through comprehensive experiments.
II Preliminaries
In this section, we present background on heterogeneity-aware FL algorithms and their approaches to handling data heterogeneity. To describe FL techniques, we consider a federated network with $K$ edge devices where each device $k$ holds its local dataset $\mathcal{D}_{k}$, FL aims to solve:
$\min_{w}F(w)=\sum_{k=1}^{K}F_{k}(w),$ (1)
where $F_{k}(w)=\frac{1}{|\mathcal{D}_{k}|}\sum_{(x,y)\in\mathcal{D}_{k}}\ell(w;x,y)$ represents the local objective function of device $k$.
FedProx [14] addresses data heterogeneity by introducing a proximal term in the local objective function, enforcing local model updates to remain close to the global model:
$F_{k}^{Prox}(w)=F_{k}(w)+\frac{\mu}{2}|w-w^{g_{e}}|^{2},$ (2)
where $g_{e}$ denotes global epoch and $\mu$ controls client drift. This proximal regularization allows FedProx to achieve more stable convergence even when data across devices is highly heterogeneous.
FedDyn [15] utilizes dynamic regularization to align local optimization objectives with the global goal by adapting to local optimization paths. The local objective in FedDyn are as follows:
$F_{k}^{Dyn}(w)=F_{k}(w)+h^{g_{e}}(w^{g_{e}})^{T}(w-w^{g_{e}})+\frac{\alpha}{2}|w-w^{g_{e}}|^{2},$ (3)
where $\alpha$ is a scaling factor and $h^{g_{e}}$ captures optimization trajectory differences. This adaptive approach allows the model to dynamically update regularization, ensuring a better fit to each device’s unique data distribution.
FedRS [16] introduces a restricted softmax method to tackle label distribution heterogeneity as follows:
$\psi_{i,c}^{k}=\frac{\exp(\alpha_{c}^{k}(w_{c}^{k})^{T}h_{i}^{k})}{\sum_{j=1}^{C}\exp(\alpha_{j}^{k}(w_{j}^{k})^{T}h_{i}^{k})},$ (4)
where $h_{i}^{k}$ is the feature vector of the $i$-th sample and $w_{c}^{k}$ is the classifier weight for class $c$ on edge device $k$, $\alpha_{c}^{k}$ is a scaling factor to limit updates of missing classes while maintaining normal softmax behavior for observed classes.
FedSAM [17] incorporates Sharpness-Aware Minimization (SAM) to enhance model generalization on heterogeneous data. SAM modifies the objective to minimize sharp local optima by perturbing gradients:
$w_{k}^{SAM}=w_{k}^{l_{e}}-\rho\frac{\nabla F_{k}(w_{k}^{l_{e}}+\epsilon)}{|\nabla F_{k}(w_{k}^{l_{e}}+\epsilon)|},$ (5)
where $l_{e}$ denotes local epoch and $\epsilon=\rho\frac{\nabla F_{k}(w_{k}^{l_{e}})}{|\nabla F_{k}(w_{k}^{l_{e}})|}$ is a perturbation that stabilizes convergence by directing models toward flatter minima.
FedSpeed [18] accelerates convergence and enhances consistency through prox-correction and gradient perturbation:
$w^{g_{e},l_{e}+1}_{k}=w^{g_{e},l_{e}}_{k}-\eta(\tilde{\nabla F}^{g_{e},l_{e}}_{k}-\hat{\nabla F}^{g_{e}}_{k}+\frac{1}{\lambda}(w^{g_{e},l_{e}}_{k}-w^{g_{e}})),$ (6)
where $\eta$ denotes local learning rate and $\tilde{\nabla F}^{g_{e},l_{e}}_{k}$ is a quasi-gradient, and $\hat{\nabla F}^{g_{e}}_{k}$ is a prox-correction term. This technique mitigates client drift while maintaining high generalization through flat minima search, leading to faster convergence with larger local intervals.
III Problem and Model
III-A Byzantine Attack and Non-IID Setting
In this paper, we consider FL environment vulnerable to both targeted and untargeted Byzantine attacks in medical imaging task as [19]. Specifically, the healthcare FL system comprises $K$ participating edge devices, including $B$ benign and $M$ compromised nodes, all connected to a central server. Each edge device $k$ maintains its private patient dataset $\mathcal{D}_{k}$ with varying sizes. To preserve patient privacy while developing a shared medical image classification model, edge devices collaborate through $G$ global epochs by sharing only model parameters with the central server.
III-B Plugin-based Byzantine-Resilient FL
In this section, we introduce a plugin-based approach that enhances FedAvg’s resilience against Byzantine attacks through feature-space analysis. At first, benign edge devices execute local learning and malicious edge devices craft byzantine attacks. For benign edge devices $b\in B$, local training aims to minimize the empirical loss:
$F_{b}(w)=\frac{1}{|\mathcal{D}_{b}|}\sum_{(x,y)\in\mathcal{D}_{b}}\ell(w;x,y),$ (7)
where $\ell(\cdot)$ is the cross-entropy loss function. The local model updates are performed via SGD with learning rate $\eta$:
$w_{b}^{l_{e}+1}=w_{b}^{l_{e}}-\eta\nabla F_{b}(w_{b}^{l_{e}};\mathcal{B}_{b}),$ (8)
where $\mathcal{B}_{b}$ is a randomly sampled mini-batch from $\mathcal{D}_{b}$.
Figure 2: Illustration of the proposed plugin’s scoring mechanism in feature space. The plugin measures pairwise cosine similarities (denoted as ‘cos’) between feature vectors extracted from model updates ($f_{1}$ through $f_{5}$) relative to the global model ($f_{g}$). Benign models (blue) exhibit high cosine similarity scores with other benign models, while poisoned models (red) show distinct patterns, enabling effective filtering of Byzantine attacks.
However, compromised edge devices $m\in M$ may perform malicious updates through either targeted or untargeted attacks. In targeted attacks, compromised devices modify their local updates which finished $L$ local training, to influence specific samples while maintaining overall performance:
$w_{m}^{g_{e},L}$ $\underset{attack}{\leftarrow}w_{m}^{g_{e},L}+\delta_{m},$ (9)
$\delta_{m}$ $=\lambda(w_{m}^{g_{e},L}-w^{g_{e}}),$ (10)
where $\lambda$ is a boosting factor designed to amplify the attack’s impact. For untargeted attacks, malicious devices inject arbitrary noise to degrade overall model performance:
$w_{m}^{g_{e}}$ $\underset{attack}{\leftarrow}\tau(w^{\prime}-w^{g_{e}}),$ (11)
$w^{\prime}$ $\sim\mathcal{N}(\bm{0},I),$ (12)
with scaling factor $\tau$ and standard Gaussian noise.
To defend against these attacks, our plugin performs Byzantine attack filtering using feature space analysis in three main steps: deviation analysis, feature mapping, and similarity-based rejection. First, we generate $N$ virtual samples $\{v_{n}\}_{n=1}^{N}$ from a standard normal distribution $\mathcal{N}(\bm{0},I)$ to serve as probe points for analyzing model behaviors. For each model pair $(w_{i}^{g_{e}},w_{j}^{g_{e}})$, we compute their deviations from the global model $w^{g_{e}}$:
$\Delta w_{i}$ $=w_{i}^{g_{e}}-w^{g_{e}}$ (13)
$\Delta w_{j}$ $=w_{j}^{g_{e}}-w^{g_{e}}$ (14)
These deviation vectors capture how each local model update differs from the current global model.
Next, we map these deviations to feature representations using a feature extractor $\mathbf{g}_{\phi}(\cdot)$, which is implemented as all layers of the base model except the final classification layer. For each model pair $(w_{i},w_{j})$ and virtual sample $v_{n}$, we compute:
$f_{i}$ $={f_{i}^{1},f_{i}^{2},...,f_{i}^{N}}=\mathbf{g}_{1:L-1}(v_{n};\Delta w_{i})$ (15)
$f_{j}$ $={f_{j}^{1},f_{j}^{2},...,f_{j}^{N}}=\mathbf{g}_{1:L-1}(v_{n};\Delta w_{j})$ (16)
where $\mathbf{g}_{1:L-1}$ denotes the feature extraction layers of the model (all layers except the final classification layer), and each $f_{i}^{n},f_{j}^{n}\in\mathbb{R}^{d}$ represents the $d$-dimensional feature vector for the $n$-th virtual sample.
After that, pairwise similarity between models is then quantified using the average cosine similarity across all virtual samples:
$s_{i,j}=\frac{1}{N}\sum_{n=1}^{N}\cos(f_{i}^{n},f_{j}^{n})$ (17)
where $\cos(f_{i}^{n},f_{j}^{n})$ measures the angular similarity between feature vectors, with higher values indicating more similar behavior patterns.
As shown in Fig. 2, benign models naturally cluster together with high similarity scores, while Byzantine models exhibit distinct patterns. For each model update $w_{k}$, we calculate its average cosine similarity with other updates:
$\bar{s}_{k}=\frac{1}{K-1}\sum_{j\neq k}s_{k,j}$ (18)
We then sort these average similarity scores $\{\bar{s}_{k}\}_{k=1}^{K}$ in ascending order and define set $\mathcal{S}$ as:
$\mathcal{S}=\{k\mid\pi(\bar{s}_{k})>M\}$ (19)
where $\pi(\bar{s}_{k})$ represents the position of $\bar{s}_{k}$ in the sorted scores, effectively excluding the $M$ updates with lowest similarity scores as potential Byzantine attacks. Subsequently, Byzantine-resilient global model is then updated through averaging of the remaining model updates:
$w^{g_{e}+1}=\frac{1}{|\mathcal{S}|}\sum_{k\in\mathcal{S}}w_{k}$ (20)
Lastly, central server broadcast the aggregated model to all edge devices for the next $g_{e}$ of local training. This process continues iteratively until the desired model performance is achieved or a maximum number of $g_{e}$ is reached. Overall procedure of proposed plugin is described in Algorithm 1.
Input: Local models $\{w_{k}^{g_{e}}\}_{k=1}^{K}$, global model $w^{g_{e}}$, number of malicious devices $M$
Output: Global model $w^{g_{e}+1}$
/* Virtual data-drive consistency scoring at central server */
1 Generate $\{v_{n}\}_{n=1}^{N}\sim\mathcal{N}(\bm{0},I)$;
2
3for $i,j\in[K]$ do
4       $f_{i}\leftarrow\mathbf{g}_{1:L-1}(v_{n};w_{i}^{g_{e}}-w^{g_{e}})$;
5       $f_{j}\leftarrow\mathbf{g}_{1:L-1}(v_{n};w_{j}^{g_{e}}-w^{g_{e}})$;
6       $s_{i,j}\leftarrow\frac{1}{N}\sum_{n=1}^{N}\cos(f_{i}^{n},f_{j}^{n})$;
7
8 end for
9
[…]
11 $\mathcal{S}=\{k\mid\pi(\bar{s}_{k})>M\}$;
12 Aggregate $w^{g_{e}+1}=\frac{1}{|\mathcal{S}|}\sum_{k\in\mathcal{S}}w_{k}$;
return $w^{g_{e}+1}$
Algorithm 1 Proposed Virtual Data-driven Byzantine Attack Filtering Plugin
IV Experiment and Results
IV-A Experiment setting
We evaluate our proposed plugin on blood cell classification task [20], implementing a ResNet-18 [21] architecture as our base model. Additionally, Gaudi 2 AI accelerator from Intel was deployed to efficiently manage the computationally demanding process, delivering scalable performance throughout our experimental work. Our experiments compare the performance of various FL techniques (FedAvg, FedProx, FedDyn, FedRS, FedSAM, FedSpeed) with and without our proposed plugin under both Targeted Model Poisoning (TMP) and Untargeted Model Poisoning (UMP) attack scenarios. To simulate realistic data heterogeneity, we distribute the data non-uniformly across $K=10$ medical edge devices following the quantity skew protocol [22]. The complete implementation details and configuration parameters are publicly available in our repository [^1] https://github.com/NAVER-INTEL-Co-Lab/gaudi-byzantine.
IV-B Results
IV-B1 Impact of Byzantine Attacks
First, we check the effect of model poisoning attacks on representative heterogeneity-aware FL techniques. Experimental results reveal the significant vulnerability of FL methods to Byzantine attacks as shown in Fig. 3. Despite handling data heterogeneity effectively, methods like FedProx, FedRS, FedDyn, FedSAM, and FedSpeed show severe performance degradation. These methods exhibit vulnerability under both targeted and untargeted attacks in our experiments. As the ratio of compromised devices increases, we observe a dramatic decrease in accuracy across all methods. Performance drops to approximately 20% when $p=0.3$, indicating their inherent weakness against malicious updates without proper defense mechanisms.
Furthermore, performance degradation varies between attack types. Untargeted attacks show notably more severe impact compared to targeted attacks across all FL methods. These methods maintain acceptable accuracy under targeted attacks with lower compromise ratios ($p=0.1$). However, they suffer catastrophic failure under untargeted attacks even at these lower ratios. This vulnerability originates from their primary focus on addressing data heterogeneity. As a result, they lack effective defenses against malicious updates in practical FL systems.
Additionally, we check conventional Byzantine-Resilient methods like Krum [23], Trimmed-Mean [24], and Fang [25] as baselines. These methods help demonstrate the relative impact of Byzantine attacks on heterogeneity-aware FL methods. The performance gap between these two categories is significant. This difference clearly emphasizes the severity of Byzantine attacks in heterogeneous FL environments.
(a) targeted attack
(b) untargeted attack
Figure 3: Performance comparison of heterogeneity-aware FL methods and Byzantine-Resilient FL methods under different attack scenarios. The x-axis represents the ratio of compromised devices ($p$), and the y-axis shows the maximum test accuracy (%).
IV-B2 Impact of Different Compromise Fractions
In this experiment, we evaluate the plugin’s effectiveness across different fractions of compromised devices. For TMP attacks, our plugin shows remarkable effectiveness across all FL methods as shown in Fig. LABEL:fig:plot2. Initially, vanilla FedAvg’s accuracy drops sharply from 80.06% to 19.47% without plugin as $p$ increases to 0.3. Yet, with plugin applied, FedAvg achieves accuracy above 89.68% even at $p=0.3$. Furthermore, similar improvements are observed in FedProx, FedDyn, and FedRS, where the plugin-enhanced versions maintain accuracy between 85-90% across all compromise ratios. Consequently, these results indicate that our plugin effectively neutralizes targeted attacks by identifying and filtering malicious updates.
Meanwhile, under UMP attacks, the plugin demonstrates strong defense capabilities despite the more challenging attack scenario. Specifically, all methods suffer severe degradation without plugin, dropping to around 17-19% accuracy at $p=0.1$. Yet, FedAvg, FedProx, FedDyn, and FedRS with plugin maintain accuracy between 65-70% even at $p=0.3$. Nevertheless, FedSAM and FedSpeed show limited improvement with plugin under UMP attacks, maintaining only 25-30% accuracy at higher compromise ratios. Indeed, such performance pattern suggests that untargeted attacks pose a greater challenge due to their random nature of perturbations.
Additionally, comparing TMP and UMP scenarios reveals interesting patterns in our plugin’s defensive capabilities. Notably, our plugin provides stronger protection against TMP attacks, with higher accuracy compared to UMP scenarios. Subsequently, FedAvg, FedProx, FedDyn, and FedRS consistently show the best plugin-enhanced performance in both attack types, maintaining above 85% accuracy in TMP and 65% in UMP at $p=0.3$. As $p$ increases, the performance gap between vanilla and plugin-enhanced versions widens, reaching maximum difference at $p=0.3$. Thus, our plugin becomes increasingly crucial as the threat level intensifies.
IV-B3 Impact of Plugin with Different Non-IID Degrees
Lastly, we check the plugin’s effectiveness improves significantly as data heterogeneity decreases ($\alpha$ increases), under TMP attacks at $p=0.3$. Notably, all vanilla methods remain at 20% accuracy regardless of $\alpha$ as shown in Fig. LABEL:fig:plot3. In contrast, plugin-enhanced versions show dramatic improvement beyond $\log\alpha=1$. FedAvg, FedProx, FedDyn, and FedRS with plugin achieve 85-90% accuracy at $\log\alpha=1$. Meanwhile, FedSAM and FedSpeed with plugin reach 70-80% accuracy. Furthermore, this trend indicates that our plugin’s defense mechanism becomes more effective usually as data distributions become more homogeneous.
Moreover, impact of data heterogeneity is more pronounced in untargeted model poisoning scenarios. Specifically, vanilla methods consistently show poor performance around 17-20% across all $\alpha$ values. Additionally, plugin-enhanced methods demonstrate gradual improvement as $\log\alpha$ increases from -1 to 2. FedAvg, FedProx, FedDyn and FedRS with plugin reach 74-75% accuracy at $\log\alpha=2$. Subsequently, FedSAM and FedSpeed follows with 64-65% accuracy. Thus, these results suggest that untargeted attacks remain challenging even with reduced data heterogeneity. Particularly, in highly heterogeneous settings ($\log\alpha<0$), even plugin-enhanced methods struggle to maintain performance above 30%.
Furthermore, performance gap between TMP and UMP scenarios reveals interesting insights about our plugin’s behavior. Consequently, at high $\log\alpha$ values, the gap between TMP and UMP performance is approximately 15-20%. As a result, this consistent difference suggests inherent limitations in defending against untargeted attacks. Nevertheless, steady improvement in both scenarios as $\alpha$ increases demonstrates our plugin’s adaptability. Therefore, while data heterogeneity affects plugin performance, it maintains significant defensive capabilities even under challenging conditions. Ultimately, these findings highlight the importance of considering data distribution characteristics when deploying Byzantine-resilient FL systems.
V Conclusion
In this paper, we propose a intuitive plugin-based approach for achieving Byzantine resilience in heterogeneity-aware FL techniques. Specifically, our solution can be applied seamlessly without modifying core FL algorithms. Through comprehensive experiments on medical image classification task, we validate our proposed approach. Indeed, our virtual data-driven consistency scoring effectively identifies and filters out malicious updates. Furthermore, the plugin maintains high accuracy while providing robust defense. As a result, we achieve significant performance improvements compared to vanilla FL methods. Thus, our plugin offers a practical solution for securing FL systems against Byzantine attacks.